# Koenigs functions in the subcritical and critical Markov branching processes with Poisson probability reproduction of particles

### v1

@@ 3.1 Integration of the reciprocal power series in the neighbourhood of $s=0$
For any values $0<\lambda\leq 1,$ the table of $B_{n,k}(p_{\bullet})$ is given by
$\left(\begin{array}[]{ccccc}(\lambda-e^{\lambda})&0&0&0&0\\ \lambda^{2}&(\lambda-e^{\lambda})^{2}&0&0&0\\ \lambda^{3}&3(\lambda-e^{\lambda})\lambda^{2}&(\lambda-e^{\lambda})^{3}&0&0\\ \lambda^{4}&4(\lambda-e^{\lambda})\lambda^{3}+3\lambda^{4}&6(1-e^{\lambda})^{2}\lambda^{2}&(\lambda-e^{\lambda})^{4}&0\\ \lambda^{5}&5(\lambda-e^{\lambda})\lambda^{4}+10\lambda^{5}&10(1-e^{\lambda})^{2}\lambda^{3}+15(1-e^{\lambda})\lambda^{4}&10(1-e^{\lambda})^{3}\lambda^{2},&(\lambda-e^{\lambda})^{5}\\ \end{array}\right)$
@@
$A(s)=\int^{s}_{x=0}\frac{(\lambda-1)dx}{e^{-\lambda}e^{\lambda s}-s}.$
@@ 4.1 In the neighborhood of the point $s=0$
Let us consider the function
$U(s)=\int^{s}_{x=0}\frac{Kdx}{f(x)}=\int^{s}_{x=0}\frac{dx}{e^{s-1}-s}.$
@@
$a_{5}=-1+2!\{5(1-e)+10\}-3!\{10(1-e)^{2}+15(1-e)\}+4!.10(1-e)^{3}-5!(1-e)^{5},$
For $\lambda=1$ the table of $B_{n,k}(p_{\bullet})$
$\left(\begin{array}[]{cccccc}(1-e)&0&0&0&0\\ 1&(1-e)^{2}&0&0&0\\ 1&3(1-e)&(1-e)^{3}&0&0&0\\ 1&4(1-e)+3&6(1-e)^{2}&(1-e)^{4}&0&0\\ 1&5(1-e)+10&10(1-e)^{2}+15(1-e)&10(1-e)^{3}&(1-e)^{5}&0\\ 1&25+6(1-e)&60(1-e)+15(1-e)^{2}+15&45(1-e)^{2}+20(1-e)^{3}&15(1-e)^{4}&(1-e)^{6}\\ \end{array}\right)$
The integration takes form
@@
$U(0)=0,\quad U^{\prime}(0)=e,\quad U^{(n)}(0)=ea_{n-1}.$
@@
Finally, we note that for them is valid that,
@@
Also, see the Example $7$, [12], (with a difference in the notation). The $U(s)$ and the $C(s)$ functions are shown in Figure 7.
Figure 7: Graphics of $U(s)$ (left) and $C(s)=e^{U(s)}$ (right).
@@ 4.2 In the neighborhood of the point $s=1$
$c_{0}=1,\quad c_{1}=\frac{-1}{3},\quad c_{2}=\frac{1}{3}\frac{1}{3!},\quad c_{3}=\frac{1}{3!}\frac{1}{15},\quad c_{4}=\frac{-1}{3!.45},\quad c_{5}=\frac{-5}{3!21.9}$

### v2

@@ 3.1 Integration of the reciprocal power series in the neighbourhood of $s=0$
Example 3.1.
For instance, the table of $B_{n,k}(p_{\bullet}),n=5$ for any values $0<\lambda\leq 1$, is given by
$\begin{pmatrix}\lambda-e^{\lambda}&0&0&0&0\\ \lambda^{2}&(\lambda-e^{\lambda})^{2}&0&0&0\\ \lambda^{3}&3(\lambda-e^{\lambda})\lambda^{2}&(\lambda-e^{\lambda})^{3}&0&0\\ \lambda^{4}&4(\lambda-e^{\lambda})\lambda^{3}+3\lambda^{4}&6(1-e^{\lambda})^{2}\lambda^{2}&(\lambda-e^{\lambda})^{4}&0\\ \lambda^{5}&5(\lambda-e^{\lambda})\lambda^{4}+10\lambda^{5}&10(1-e^{\lambda})^{2}\lambda^{3}+15(1-e^{\lambda})\lambda^{4}&10(1-e^{\lambda})^{3}\lambda^{2}&(\lambda-e^{\lambda})^{5}\end{pmatrix}$ (10)
@@
$A(s)=\int^{s}_{x=0}\frac{(\lambda-1)dx}{e^{-\lambda}e^{\lambda x}-x}.$
@@ 4.1 In the neighborhood of the point $s=0$
Let us consider the function $U(s),U(0)=0,$ defined by the integral for $\lambda=1$ as
$U(s)=\int^{s}_{x=0}\frac{Kdx}{f(x)}=\int^{s}_{x=0}\frac{dx}{e^{x-1}-x},\quad U^{\prime}(s)=\frac{1}{e^{s-1}-s}\quad U^{\prime}(0)=\frac{1}{e^{-1}}=e.$
@@
$a_{5}=-1+2!\{5(1-e)+10\}-3!\{10(1-e)^{2}+15(1-e)\}+4!10(1-e)^{3}-5!(1-e)^{5}.$
Following the previous Example 3.1, the tabulated values of $B_{n,k}(p_{\bullet})$ can be obtained after direct substitution of $\lambda=1$ in the matrix (10).
@@
$U(0)=0,\quad U^{\prime}(0)=e,\quad U^{(n)}(0)=ea_{n-1},\quad n=2,3,.$...
@@
Finally, we note that for them is valid that, $C(0)=1,C^{\prime}(0)=e$, and
@@
Also, see the Example $7$, [12], (with a difference in the notation).
Remark 1.
Remamber the functions
$A(s)=\int^{s}_{x=0}\frac{(\lambda-1)dx}{e^{-\lambda}e^{\lambda x}-x},\quad U(s)=\int^{s}_{x=0}\frac{dx}{e^{x-1}-x}.$
Denote the function $U_{\lambda}(s)=\frac{A(s)}{\lambda-1}$ as follows,
$U_{\lambda}(s)=\frac{A(s)}{\lambda-1},\quad U_{\lambda}(s)=\int^{s}_{x=0}\frac{dx}{e^{-\lambda}e^{\lambda x}-x},$
Then for the critical case, the function $U_{1}(s):=U(s)$.
The graphics of the functions $U_{\lambda}(s)$ and $U(s)$ and respectively $C_{\lambda}(s)=e^{U_{\lambda}(s)}$ and $C(s)=e^{U(s)}$ are shown in Figure 7.
Figure 7: Graphics of $U_{\lambda}(s)$ and $U(s),U(0)=0,U^{\prime}(0)=e$ (left) and respectively $C_{\lambda}(s)=e^{U_{\lambda}(s)}$ and $C(s)=e^{U(s)},C(0)=1$ (right). The dashed line on right figure points out the minimal value at $C(0)=1$.
@@ 4.2 In the neighborhood of the point $s=1$
$c_{0}=1,\quad c_{1}=\frac{-1}{3},\quad c_{2}=\frac{1}{3}\frac{1}{3!},\quad c_{3}=\frac{1}{3!}\frac{1}{15},\quad c_{4}=\frac{-1}{3!45},\quad c_{5}=\frac{-5}{3!21}\frac{1}{9},$...